# GR3EN: Generative Relighting for 3D Environments

### v1

@@ 2 Related Work
Inverse rendering. Early solutions for scene relighting were built upon inverse rendering systems, which decompose a scene’s appearance into its constituent factors of materials, illumination, and geometry (or some subset of these, assuming the others are known) [7, 30, 41, 31]. These methods either directly invert a simple image formation model [15], or iteratively optimize these factors [41, 5, 10, 20, 19], typically using simplified models for reflectance and lighting. Early attempts at inverse rendering parameterize scene geometry using meshes or depth maps, while the rise of radiance fields [25, 18] enabled effective approaches for inverse rendering that parameterize a scene using the weights of a neural network [44, 32, 4, 17, 24] or a set of 3D Gaussians [12]. While this has substantially improved inverse rendering results by enabling gradient-based optimization through a differentiable image formation process, this line of work is still limited by the central challenge of inverse rendering: that it is highly ill-posed. As such, these methods rely on simplifying assumptions of the scene or brittle priors over scene content, which severely limits photorealism, or rely on an abundance of input data (e.g., images of both novel views and novel lighting conditions), which limits the settings in which these models can be used.
2D image relighting. Many recent approaches have attempted to circumvent the challenging inverse rendering decomposition problem by directly learning to relight images in a fully data-driven manner. These approaches have trained feed-forward neural networks [27, 33], or generative image models such as Generative Adversarial Networks (GANs) [36, 26, 34, 8] and more recently diffusion models [23, 43, 16, 38, 39, 42] to generate plausible relit images, conditioned on a desired target lighting. While these methods have been effective at producing plausible 2D results, they do not produce 3D outputs that can be used for view synthesis or other graphics applications.
Video relighting. Building on recent advancements in video diffusion models, a number of works have addressed the challenge of 2D video relighting by using diffusion models [46, 11, 9, 13, 21]. However, these approaches provide limited control over lighting, only allowing editing coarse control such as modifying directional lights or using text prompts to describe the target light.
@@ 5.2 Baselines
Conditioning for DiffusionRenderer. DiffusionRenderer relights scenes based on an environment map. Therefore for each target illumination we use our relit NeRF to render a target environment map. Since DiffusionRenderer is limited to at most 57 frames, we evaluate it on the first 57 frames of our elliptical camera path and report its results on this subset.
@@ 5.3 Real-World Scene Relighting Results
We evaluate our method on captured scenes from the Eyeful Tower dataset [40]. This datasets consists of high-quality captures of indoor environments. We show qualitative results of our entire pipeline (reconstruction, relighting, and 3D distillation) in Figures 1, 2, 4, 5, and 9. Perhaps surprisingly, our model was never trained on light fixture shapes present in this data, such as the large rectangular ceiling light sources in the top half of Figure 4, but it successfully generalizes to out-of-distribution lighting conditions.
@@
Figures 10 and 11 show the results of our relighting model, demonstrating that our diffusion model provides fine-grained control over the illumination of the scene, rendering realistic global illumination effects like cast shadows and reflections.
@@
User study. We conduct a user study comparing our method with LightLab [23], DiffusionRenderer [21], and the original input video (which has no artifacts but does not respect target lighting). Users were shown pairs of videos with identical camera paths and a text description of the target lighting condition, then asked to evaluate frame-to-frame consistency, object identity preservation, and lighting similarity.
@@ 5.4 Synthetic Scene Relighting Results
| PSNR | SSIM | LPIPS | |
| Channel concatenation | 17.060 | 0.648 | 0.172 |
| w.o. high noise bias | 23.823 | 0.637 | 0.192 |
| Ours | 24.765 | 0.741 | 0.130 |
| (b) All indoor lights turned off, exterior light only | |||||
| (c) Partial ceiling lamps turned on, exterior light disabled | |||||
| (d) Selected ceiling lamps on with varying emission colors | |||||
In Table 3 we report our relighting results compared with DiffusionRenderer (Cosmos version) [21], LightLab [23] and Light-A-Video [46]. Although DiffusionRenderer can theoretically relight 3D scenes, its control over light sources is extremely limited. As shown in Figure 12, DiffusionRenderer fails to accurately modify visible light sources as well as exterior light sources that are not directly visible. While LightLab can edit the color of the lights, its control is limited, and it fails to turn on exterior light sources.
@@ Acknowledgment
We would like to thank Gene Chou, Rundi Wu, Ruiqi Gao, Songyou Peng and Alexander Raistrick for their valuable help with implementation. We also thank Amir Hertz and Eric Tabellion for providing LightLab assets. We appreciate the insightful discussions with Connie He, Georgios Kopanas, Janne Kontkanen, Michael Okunev, Ben Poole, and Rick Szeliski. Finally, we thank all participants who took part in our user study.
| Input | LightLab | Light-A-Video | DiffusionRenderer | Ours | Ground Truth |
Supplementary Material
@@ Appendix B Real-world 3D Environment Relighting
We conduct experiments on four scenes from the Eyeful Tower dataset [40]: office1a, office_view2, kitchen, and seating_area. For each scene, we first mask the most prominent light sources in 3D. This is done by selecting image pixels in 2D and back-projecting them to 3D, and modifying the color of the 3D points corresponding to the light sources to and other points to . The resulting 3D light masks are then rendered along the same elliptical camera path to obtain per-frame 2D light masks that are aligned with the input video frames.
@@ B.1 Details of User Study
Figure 13 shows the interface of our user study. Users were asked to compare the two videos and determine which better satisfies three criteria: frame-to-frame consistency, object identity preservation, and lighting similarity to the target condition. Each comparison was presented in randomized order to mitigate position bias, and responses were aggregated across all raters. The lower win rate against Input (68.5%) is expected, as the original video naturally preserves temporal consistency and object identity perfectly—yet users still preferred our method, indicating they value the improved lighting fidelity our approach provides.

### v2

@@ 2 Related Work
Inverse rendering. Early solutions for scene relighting were built upon inverse rendering systems, which decompose a scene’s appearance into its constituent factors of materials, illumination, and geometry (or some subset of these, assuming the others are known) [7, 39, 51, 40]. These methods either directly invert a simple image formation model [20], or iteratively optimize these factors [51, 5, 10, 26, 25, 53, 11], typically using simplified models for reflectance and lighting. Early attempts at inverse rendering parameterize scene geometry using meshes or depth maps, while the rise of radiance fields [34, 23] enabled effective approaches for inverse rendering that parameterize a scene using the weights of a neural network [56, 41, 4, 22, 31, 59] or a set of 3D Gaussians [15, 13]. While this has substantially improved inverse rendering results by enabling gradient-based optimization through a differentiable image formation process, this line of work is still limited by the central challenge of inverse rendering: that it is highly ill-posed. As such, these methods rely on simplifying assumptions of the scene or brittle priors over scene content, which severely limits photorealism, or rely on an abundance of input data (e.g., images of both novel views and novel lighting conditions), which limits the settings in which these models can be used.
2D image relighting. Many recent approaches have attempted to circumvent the challenging inverse rendering decomposition problem by directly learning to relight images in a fully data-driven manner. These approaches have trained feed-forward neural networks [36, 42, 16, 55, 33], or generative image models such as Generative Adversarial Networks (GANs) [46, 35, 43, 8] and more recently diffusion models [30, 54, 21, 48, 49, 52, 45, 18, 24, 12] to generate plausible relit images, conditioned on a desired target lighting. While these methods have been effective at producing plausible 2D results, they do not produce 3D outputs that can be used for view synthesis or other graphics applications. A concurrent work, LuxRemix [28], proposes a relighting framework based on image diffusion models that first decomposes a scene into its constituent light sources, followed by a trained multi-view harmonization model to propagate lighting across different views for scene relighting. In contrast, our method directly achieves consistent room-scale relighting without requiring explicit light source decomposition or view-wise harmonization.
Video relighting. Building on recent advancements in video diffusion models, a number of works have addressed the challenge of 2D video relighting by using diffusion models [58, 14, 9, 17, 27, 32]. However, these approaches provide limited control over lighting, only allowing editing coarse control such as modifying directional lights or using text prompts to describe the target light.
@@ 3.3 Reconstruction and 3D Distillation
Throughout the main paper, our results are obtained by using Zip-NeRF for reconstruction and 3D distillation due to its high quality [6]. However, nothing about our approach is specific to any particular radiance field technique, and one could instead use a real-time (but slightly lower quality) approach such as 3D Gaussian Splatting [23].
@@ 5.1 Implementation Details
We plan to release our code and trained relighting model weights upon acceptance of our paper.
@@ 5.2 Baselines
Conditioning for DiffusionRenderer. DiffusionRenderer relights scenes based on an environment map. Therefore for each target illumination we use our relit NeRF to render a target environment map. Since DiffusionRenderer supports at most 57 frames, we evaluate it on the first 57 frames of our camera path and report its results on this subset.
@@ 5.3 Real-World Scene Relighting Results
We evaluate our method on captured scenes from the Eyeful Tower dataset [50]. This datasets consists of high-quality captures of indoor environments. We show qualitative results of our entire pipeline (reconstruction, relighting, and 3D distillation) in Figures 1, 2, 4, and 5. Perhaps surprisingly, our model was never trained on light fixture shapes present in this data, such as the large rectangular ceiling light sources in the top half of Figure 4, but it successfully generalizes to out-of-distribution lighting conditions.
@@
Figures 8 and 9 show the results of our relighting model, demonstrating that our diffusion model provides fine-grained control over the illumination of the scene, rendering realistic global illumination effects like cast shadows and reflections.
@@
User study. We conducted a user study comparing our method with LightLab [30], DiffusionRenderer [27], and the original input video (which has high quality but does not respect the target lighting). Users were shown pairs of videos with identical camera paths and a text description of the target lighting condition, then asked to evaluate frame-to-frame consistency, object identity preservation, and lighting similarity.
@@ 5.4 Synthetic Scene Relighting Results
In Table 3 we report our relighting results compared with DiffusionRenderer (Cosmos version) [27], LightLab [30] and Light-A-Video (WAN 2.1 version) [58]. Although DiffusionRenderer can theoretically relight 3D scenes, its control over light sources is extremely limited. As shown in Figure 10, DiffusionRenderer fails to accurately modify visible light sources as well as exterior light sources that are not directly visible. While LightLab can edit the color of the lights, its control is limited, and it fails to turn on exterior light sources.
@@ 5.6 Additional Applications
| PSNR | SSIM | LPIPS | |
| Channel concatenation | 17.06 | 0.648 | 0.172 |
| w.o. high noise bias | 23.82 | 0.637 | 0.192 |
| Ours | 24.77 | 0.741 | 0.130 |
@@ Acknowledgment
We would like to thank Gene Chou, Rundi Wu, Ruiqi Gao, Songyou Peng and Alexander Raistrick for their valuable help and suggestion with implementation. We also thank Amir Hertz and Eric Tabellion for providing LightLab assets. We appreciate the insightful discussions with Connie He, Aleksander Holynski, Georgios Kopanas, Janne Kontkanen, Michael Okunev, Ben Poole, Rick Szeliski, and Dimitris Tzionas. Finally, we thank all participants who took part in our user study.
Appendix
@@ Appendix B Real-world 3D Environment Relighting
We conduct experiments on four scenes from the Eyeful Tower dataset [50]: office1a, office_view2, kitchen, and seating_area. For each scene, we first mark the most prominent light sources in 3D. This is done by selecting image pixels in 2D and back-projecting them to 3D, and modifying the color of the 3D points corresponding to the light sources to and other points to . The resulting 3D light masks are then rendered along the same elliptical camera path to obtain per-frame 2D light masks that are aligned with the input video frames.
@@ B.1 Details of User Study
Figure 11 shows the interface used in our user study. Each of 12 users is asked to make 18 comparisons, making up a total of 216 responses. Users were asked to compare the two videos and determine which better satisfies three criteria: frame-to-frame consistency, object identity preservation, and lighting similarity to the target condition. Each comparison was presented in randomized order to mitigate position bias, and responses were aggregated across all raters. The lower win rate against Input (68.5%) is expected, as the original video naturally preserves temporal consistency and object identity perfectly—yet users still preferred our method, indicating they value the improved lighting fidelity our approach provides.
@@ Appendix C Data Generation Pipeline Visualization
| (b) All indoor lights turned off, exterior light only | |||||
| (c) Partial ceiling lamps turned on, exterior light disabled | |||||
| (d) Selected ceiling lamps on with varying emission colors | |||||